Find the equilibrium level of income for the two economies (A & B) that has the following characteristics
A: C = 1000 + 0.8Y and I = 800
B: C = 1000 + 0.8Y and I = 500
From (i) above can we conclude that having the same or similar consumption function would always yield similar equilibrium level of income? Provide reasons for your answer.
What accounts for the difference in the equilibrium level of income for both countries
1
Expert's answer
2020-11-02T06:58:02-0500

the equilibrium level of income for the A:

Y=C+I

Y=1000+0.8Y+600

Y-0.8Y=1600

0.2Y=1600

Y=8000


the equilibrium level of income for the B:

Y=C+I

Y=1000+0.8Y+500

Y-0.8Y=1500

0.2Y=1500

Y=7500

We can say that this is because the "national income—total expenditure" or "income—expenditure" equilibrium model»

The difference in the equilibrium level of income for both countries isdue to the level of investment in the country's economy